May 1 Butler invested $20,000 cash and his tugboat valued at $90,000 in the business in...

May 1

Butler invested $20,000 cash and his tugboat valued at $90,000 in the business in exchange for common Stock

May 2

Butler paid $3,000 cash for office equipment to help him keep track of business activities.

May 3

Butler bought boating supplies costing $2,500 on credit.

May 4

Butler paid the river master $500 cash for the first month's dock rental.

May 5

Butler paid $1,800 cash for a six-month insurance policy.

May 10

Butler received $2,000 cash from clients for his first tour

May 12

Butler provided a $3,500 tour on credit, the customer has agreed to pay within 10 days.

May 19

Butler paid for the boating supplies originally purchased on May 3.

May 22

Butler receives payment on the account from the client entry on May 12.

May 31

Butler paid his crew member a salary of $1,000

May 31

The company paid $2,000 cash in dividends to the owner. (sole stockholder)


Need help making a trial balance!
Answer #1
Trial Balance
For the Month Ended May 31
Debit Credit
Cash $    14,700
Tugboat $    90,000
Common Stock $ 110,000
Office Equipment $      3,000
Supplies $      2,500
Accounts Payable $            -  
Rent Expense $         500
Prepaid Insurance $      1,500
Insurance Expense $         300
Revenue $      5,500
Accounts Receivable $            -  
Salary Expense $      1,000
Dividend/Drawings $      2,000
Total $ 115,500 $ 115,500

Here is the formula version

Trial Balance
For the Month Ended May 31
Debit Credit
Cash =20000-3000-500-1800+2000-2500+3500-1000-2000
Tugboat 90000
Common Stock =20000+90000
Office Equipment 3000
Supplies 2500
Accounts Payable =2500-2500
Rent Expense 500
Prepaid Insurance =1800-300
Insurance Expense 300
Revenue =2000+3500
Accounts Receivable =3500-3500
Salary Expense 1000
Dividend/Drawings 2000
Total =SUM(D6:D18) =SUM(E6:E18)